OpenStudy (clamin):

MEDAL!! PLEASE HELP!! 7x^2 + 13x = 2 how do i solve this??

pooja195 (pooja195):

Put it into standered form first \[\huge~\rm~ \bf ax^2+bx+c\]

pooja195 (pooja195):

Any ideas on how we would do this? (hint move the 2)

pooja195 (pooja195):

@clamin

Zeronknight (zeronknight):

@clamin Are you there? :)

OpenStudy (clamin):

hold on

OpenStudy (clamin):

whats my first step gonna be??

OpenStudy (anonymous):

combine like terms

jagr2713 (jagr2713):

Well can this be factored @clamin

OpenStudy (triciaal):

@megan.brennan there are no like terms to combine

jagr2713 (jagr2713):

Its a question @clamin

OpenStudy (anonymous):

@triciaal Why not?

jagr2713 (jagr2713):

Because x and x^2 are not the same :/

OpenStudy (triciaal):

x^2 is not the same as x

OpenStudy (triciaal):

like apples and bananas are both fruits but not the same fruit

OpenStudy (clamin):

do i factor this?

OpenStudy (triciaal):

set equal to zero first yes

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

(x - 2) • (7x + 1)

OpenStudy (clamin):

theres no number that goes with 13

jagr2713 (jagr2713):

Ok, so we have to use what formula?

OpenStudy (anonymous):

Now set them to zero and solve for x

OpenStudy (clamin):

i think i got it

OpenStudy (clamin):

1/7, -2

OpenStudy (anonymous):

almost

jagr2713 (jagr2713):

Correct :D

OpenStudy (anonymous):

x - 2=0 7x + 1=0

OpenStudy (anonymous):

x=2 x=-1/7

jagr2713 (jagr2713):

Its -2 @mthompson440

OpenStudy (triciaal):